Junior Cycle · Probability ⏱ Tested in 1 exam question since 2023 Lesson 6 of 7

JCHL Probability 6: Miscellaneous

This tutorial is for members

Read the lesson notes free below — subscribe to watch the full worked video and download the PDF.

Subscribe for full access · €199/yr

You’ll learn

  • P(event happens) = 1 − P(event does NOT happen)
  • 1 stands for all the possible outcomes (100%); subtract the part you do not want to get the part you do want
  • This shortcut is very useful in harder "at least" type questions

Formula used

probability it DOES happen = 1 − probability it does NOT happen

Most recent exam questions

  • 2023 Q5(c)

💬 Stuck on something in this lesson?

Send your question with this lesson’s name attached — you’ll get a reply by email.